40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Message ID: 1910     Entry time: Wed May 19 09:25:34 2021
Author: Paco 
Type: Noise Hunting 
Category: DOPO 
Subject: 316 Hz noise 

[Paco]

- Have been investigating 316 Hz noise in the control signal for the DOPO lock. Here is a list of some things that have been ruled out, mostly electrical:

  - EOM power supply --> noise still present in DOPO transmission
  - RFPD DC out --> no funky ground loops with scope (also looking at demod signal in different channel), noise still visible in transmission
  - RFPD power supply --> noise still visible in transmission...
  - Pump laser intensity (upstream pickoff) --> not a great test because pickoff optics are also on the optical table..
  - 2 x SR560s --> No effect after bypassing
  - Marconi --> same result as with anything in the loop after RFPD demod

- Things left to rule out:
  
  - Fume hood exhaust fan ** highly suspected, my phone's own cheap-o microphone power spectrum shows peaks at 316.5 Hz (!) when near the exhaust fan
  - NPRO temp controller fan --> phone audio spectrum shows line noise (60 Hz) mostly, and also 188 Hz... need to test further independently of the fume hood...

In ruling out the 6-axis translation mount on the DOPO cavity, I removed the PPKTP crystal + oven temporarily but still saw the noise. Since the resonator was no longer stable without the crystal, I needed to bring the mirrors closer and realign the output coupler from scratch. 

Restored DOPO cavity with crystal, alignment. MM efficiency ~ 35%... still optimizable.

ELOG V3.1.3-